OpenStudy (anonymous):

Give the formula for the following: 1. The perimeter of a rectangle is 52. Its length is2 more than twice the measure of its width. Find its dimensions. 2. If the diameter of a circle is the same as the radius of the circumference of the larger circle to the circumference of the smaller.

OpenStudy (anonymous):

2) the ratio is 2 1)length=17.6 breadth=8.67

OpenStudy (anonymous):

sail can you tell me how you got it

OpenStudy (anonymous):

you can just give me the formula for it

OpenStudy (anonymous):

can you tell me how sail1234

OpenStudy (anonymous):

2)cirmuference formula is 2*pie*r or pie*d ratio of ciucumof large circle/circum of small circle =2*pie*r1/2*pie*r2 =for small circle d2=r1 i.e 2*r2=r1

OpenStudy (anonymous):

how about number 1

OpenStudy (anonymous):

i alredy told wen u posted prevously see once

OpenStudy (anonymous):

what post can you repeat pls pls pls